OpenStudy (anonymous):

No idea how to do this! (x-8)(x-k) = x^2 - 5kx +m In the equation above, k and m are constants. If the equation is true for all values of x, what is the value of m? ((I'm an International student and I've never been asked a question like this..))

OpenStudy (anonymous):

\[(x-8)(x-k) = x ^{2} - 5kx + m\]

OpenStudy (anonymous):

true for all \(x\)? then try \(x=0\)

OpenStudy (anonymous):

How does saying "true for all x" denote that x = 0? My apologies, I just haven't done this is my own school before!

OpenStudy (anonymous):

it says "true for all \(x\)" apparently, so what i am saying is that if it is true no matter what \(x\) is, then it must be true if \(x=8\)

OpenStudy (anonymous):

if you replace \(x\) by \(8\) you get \[64-40k+m=0\] as one equation or \[40k-m=64\] then pick some other value for \(x\) get another equation with in \(k\) and \(m\) and solve

OpenStudy (anonymous):

it doesn't really matter what value of \(x\) you pick, pick any two and solve

hero (hero):

I figured out what m equals

hero (hero):

I think it is wise to let x equal 8 then create the first equation. Afterwards, let x = k, then create the second equation.

OpenStudy (anonymous):

@Hero @satellite73 Thank you both; just another question: So why pick 8? Could you pick let's say 6, or 4? And what "type" of Math is this? (ie: Advanced Algebra, Geometry, Calculus etc..)? Thanks

hero (hero):

You use x = 8 and x = k because you can simplify the equations better that way

hero (hero):

For example on the left side: When you use x = 8, the left side reduces to 0: When you use x = k, the left side also reduces to 0: This makes solving for m more convenient.

hero (hero):

@Dj2k12

OpenStudy (anonymous):

@Hero Ah, okay. I'm going to try and work this out, and I'll publish what I end up with here. Thanks for your help!

OpenStudy (anonymous):

\[(x-8)(x-k)=x ^{2}-5kx+m. \] So from here, I plug "8" in on the left side. \[(8-8)(8-k) = 0\] Now I have to plug 8 into "x" on the other side too, so: \[x ^{2} - 5k(8) +m = 0\] Or.. \[64-40k+m=0\] So how do I get rid of k here? Am I right so far?

OpenStudy (anonymous):

@Hero

hero (hero):

It looks right. Now do it again except replace x with k

OpenStudy (anonymous):

So if I replace k with x, and x = 8, then do I just replace k with 8? Because when I do that I get -256. @Hero

hero (hero):

Just do the same thing you did the first time but replace x with k ONLY

hero (hero):

you don't replace k with 8

OpenStudy (anonymous):

I'm not entirely getting it. I don't have anything to represent k so how can I replace x with it? @Hero. Sorry, as I said, we don't cover this type of Math in Ireland lol

hero (hero):

What is so difficult about replacing x with k everywhere you see x?

hero (hero):

pretend that k is a number and replace it with x the same way you did 8 the first time.

OpenStudy (anonymous):

Erm, plugging in isn't difficult, it's understand why and when that's a challenge. Regardless, I'll give it a shot: \[x ^{2}-5kx+m\] Now I'm replacing x with k. \[k ^{2}-5k(k) + m = 0\] \[k^2 - 5k^2 +m= 0\] \[-4k^2+m=0\] .. Now what do I do?

OpenStudy (anonymous):

@Hero

hero (hero):

add \(4k^2\) to both sides

OpenStudy (anonymous):

\[m= +4k^2?\]

hero (hero):

Yes, now set m = m and solve for k

hero (hero):

After you find k, plug the value you find back in to find m. (Systems of Equations)

OpenStudy (anonymous):

Okay, so firstly, I'm plugging my new value for M into the equation to solve for K: \[k^2 - 5k^2 + 4k^2 = 0\] And now I have 0=0.

hero (hero):

You're ignoring the first equation you created when you had \(m = 40k - 64\)

hero (hero):

I stated to set m = m, then after doing that, solve for k.

OpenStudy (anonymous):

But that's because I found out that m = 4k^2?

OpenStudy (anonymous):

Oh I see what you mean, hold up

hero (hero):

smh

OpenStudy (anonymous):

so \[4k^2 = 40k-64\] Divide across by k \[4k= 40-64\] 4k= 24 and thus, k = 6. Right so far? @Hero

hero (hero):

Never ever divide by the variable you are solving for

OpenStudy (anonymous):

@Hero Sorry, I'm just not getting this question. Thanks anyway.

hero (hero):

We're almost done. Why are you going to give up now?

hero (hero):

Clearly we have a quadratic equation left to solve.

OpenStudy (anonymous):

Because I'm utterly lost. But I guess I can keep trying. Was I right with k= 4?

OpenStudy (anonymous):

Oops, I mean 6. ^^

hero (hero):

You're not lost. You made it this far without asking for much help.

OpenStudy (anonymous):

I'm just confused because first m= 40k-64, and then it equals \[4k^2\] and I finally get k = 6 and I'm not sure where to go from here, and why I am still ignoring the first part of the equation. Nope, apparently I did not.. but I'm guess that I should bring the 4k^2 over the equals and use the -b formula?

OpenStudy (anonymous):

My first instinct would be to divide across by x so that I could have it on the left side and allow it to equal to an integer on the right; but you said to "never divide by the variable you're trying to solve" so my next guess would be to find the factors of 64 and factorize the equation, or else use the -B formula. Sorry, but as you can see, Math isn't one of my strengths.

hero (hero):

Okay, the correct quadratic equation should be \[4k^2 - 40k + 64 = 0\]

hero (hero):

Writing it this way, there's no way you can simply divide both sides by k

hero (hero):

Dividing by k eliminates one of the solution values. You do not want to eliminate solution values.

OpenStudy (anonymous):

Of course I have, but we tend to use the -B formula to sort out all Quadratic equations by using the formula \[ax^2 + bx+ c\] and plugging the numbers into the minus B formula. Regardless, I could do the same here by allowing a = 4 b= 40 and c = 64. Or else I could say (4k+16)(k-4) or something like that.. I'm not entirely certain @Hero

OpenStudy (anonymous):

Oh wait, I factorized wrongly. So -B would be my last resort.

hero (hero):

That -b formula you're referring to is called the Quadratic Formula by the way and b = -40

hero (hero):

But anyway, you do not have to use the Quadratic Formula if the quadratic equation is factorable.

OpenStudy (anonymous):

Listen, thanks for your help but I really am lost in this question. I'm going to try and perfect the Medium and Easy questions and perhaps skip the questions like these on the SAT. But again, thanks for your help up until this point but I'm just not getting this. @Hero

hero (hero):

\[4k^2 - 40k + 64 = 0 \\4(k^2 - 10k + 16) = 0 \\k^2 - 10k + 16 = 0 \\k^2 - 8k - 2k + 16 = 0 \\k(k - 8) -2(k - 8) = 0 \\(k - 8)(k - 2) = 0 \\k = 2 \\ k = 8 \]

hero (hero):

Using k = 2, you get m = 16

OpenStudy (anonymous):

@Hero. Thank you; I'll write this down and I'll try to maybe work it out on paper until it makes sense. Again, thanks for your help and patience.

hero (hero):

Well unfortunately it is not as simple as that. You'll probably get more confused before you understand it. I think you should take satellite's advice and try different values of x. It's kinda risky though.